OpenStudy (anonymous):

What is the standard form equation of the line passing through (1, –5) with a slope of –1?

OpenStudy (tkhunny):

What form do you know? Maybe start with the point-slope?

OpenStudy (anonymous):

what is the point slope form?

OpenStudy (tkhunny):

I'll take that as a "no". This is not good at all. You are given a point and the slope. Why would you not be given the Point-Slope form of a line FIRST?! What forms do you know?

OpenStudy (anonymous):

Is the point slope form y=mx+b? thats all i know now. and i know how to find the slope. which im already given

OpenStudy (tkhunny):

y = mx + b This is the Slope-Intercept form. This is very useful, but it is not quite what we need for this one. Let's see if we can figure it out. Point:(1,-5) Slope: -1 Can we use this information to find the y-intercept? y-intercept: (0,What)

OpenStudy (anonymous):

We cant use that to find that, ahh so the point slope form is y-y^1 = m(x-x^1)

OpenStudy (tkhunny):

That is it, but subscripts, of course. Slope: m = -1 Point: (1,-5) gives x1 = 1 and y1 = -5 Fill in the blanks.

OpenStudy (tkhunny):

We could also have simply substituted. Given y = mx + b, We have m = -1, x = 1, and y = -5, thus: -5 = (-1)(1) + b and we can solve for b. You need Standard Form. You wouldn't happen to know that one, would you?

OpenStudy (anonymous):

-4=b, correct? I dont know what standard form is but where did you get that 1 in (-1)(1) from?

OpenStudy (anonymous):

ax + by = c is standard form?

OpenStudy (tkhunny):

(-1)(1) are the substitutions of m = -1 and x = 1 into the term "mx". With b = -4, we have y = -x - 4. Standard form is x + y = -4. We just need to learn to recognize the various forms. We also need to get our algebra up to speed so we can fluently change between the various forms.

OpenStudy (anonymous):

Yes, I know. I took algebra 1 a few years back and now im taking algebra 2 so I'm obviously a little rusty now.

OpenStudy (anonymous):

But thank you so much for your help it refreshed my memory a bit.

OpenStudy (tkhunny):

Very good. It does seem to be coming back, even in this brief conversation. I think you'll be okay. Just keep getting up to speed.